Consider that a baseball which is made of a redwood has a mass of 7.0 kg and a specific gravity of 0.65. Then, answer the following questions. (The density of water is 10^3 kg/m3)

a) The redwood ball is tied by a thin string to the bottom of a tank of water in order to have the ball total immersed. What is the tension in the string?

b) The redwood ball is tired with a thing string. But, at this time, we put it into a tank of water in order to have the ball floated. What is the tension in the string/
Answer #1

here,

a) for the volume of the redwood, V = mass/density

V = 7/(650) = 0.01077 m^3

let the tension in the string is T

balancing the net forces

T + weight - Buyoant force = 0

T + 7 * 9.8 - 0.01077 * 1000 * 9.8 = 0

solving for T

T = 37 N

the tension in the string is 37 N

b) if the redwood is floating

the string tension may be Zero